can you help me solve the problem by factoring and using the principle of zero products (x - 2)(x - 7) = -6 I need to see the steps to understand the answer
Answer by graphmatics(170)

We will try to use the quadratic factor procedure so expand the left side of the equation:
(x-2)(x-7) = -6
x^2 -7*x -2*x + 2*7 = -6
x^2 - 9*x + 14 = -6
x^2 - 9*x + 14 + 6 = 0
x^2 - 9*x + 20 = 0
(x-5)(x-4) = 0
So:
x-5 = 0 and x-4 = 0
x = 5 and x = 4